OpenStudy (anonymous):

Let f(x) = x - 6 and g(x) = x^2 - x + 4. Find g(f(x)). Question 10 options: 1) x2 - 13x + 46 2) x2 - 11x + 34 3) x2 - 2x - 2 4) x2 - x - 2

OpenStudy (anonymous):

g(x-6)=(x-6)^2-(x-6)+4 =x^2+36-12x-x+6+4 =x^2-13x+46